# Towards Data-Driven Adaptive Exoskeleton Assistance for Post-stroke Gait

### v1

@@ II-C1 Unimpaired Dataset
The Unimpaired Dataset was collected from individuals without neuromotor gait impairment (6 participants, 4 M / 2 F, age 29±3.929$\pm$ 3.929 ± 3.9 yrs). We sensorized their left leg, as all participants post-stroke were left-paretic. As shown in Figure 3 on the left, all the IMUs were laterally positioned on the participants’ left leg, aligned with the sagittal plane. Participants walked on the instrumented treadmill for 6 trials of 4 minutes each. Each trial featured randomized speed conditions ranging from 0.6 m/s to 1.4 m/s. For the first 2 minutes, the treadmill maintained a constant speed of 1 m/s before ramping up to either 1.2 m/s or 1.4 m/s, or down to 0.8 m/s or 0.6 m/s for the final 2 minutes. Half of the participants started with acceleration, while the others began with deceleration. The speed variations were implemented to capture a range of walking patterns and enhance dataset diversity. On average, 2118±4422118$\pm$ 4422118 ± 442 strides were collected per participant, resulting in a total of 127101271012710 strides.
@@ II-C2 Post-stroke Dataset
The Post-stroke Dataset was collected from 4 left-paretic stroke survivors (4 M, age $\pm 5.5$ 5.5 yrs). Participants wore the ankle robot on their paretic limb. Each participant completed 6 trials on the instrumented treadmill, comprising of two 2 min trials for each a self-selected comfortable walking speed (CWS: 0.65±0.130.65$\pm$ 0.130.65 ± 0.13 m/s), a slow walking speed (SWS: 0.5±0.080.5$\pm$ 0.080.5 ± 0.08 m/s), and a fast walking speed (FWS: 0.8±0.180.8$\pm$ 0.180.8 ± 0.18 m/s). Participants post-stroke demonstrated high variability in their self-selected CWS, which ranged from 0.5 to 0.8 m/s. On average, each participant performed 334±79334$\pm$ 79334 ± 79 strides, which resulted in a total of 133413341334 strides.
@@ III-A Feature Importance
We conducted a feature importance analysis to reduce the number of streamed channels. The results are summarized in Table I. Randomizing the acceleration measurements increased RMSE by +0.04±0.03+0.04$\pm$ 0.03+ 0.04 ± 0.03 Nm/kg, the MAE by +0.03±0.02+0.03$\pm$ 0.02+ 0.03 ± 0.02 Nm/kg, and decreased the $\mathrm{R}^{2}$ score by 0.09±0.05-0.09$\pm$ 0.05- 0.09 ± 0.05 on average. Despite accelerations being integral to inverse dynamics, their omission affected prediction accuracy the least, compared to orientation and gyroscope data. We excluded the accelerations from our subsequent model training to ensure the targeted 100 Hz throughput on the microcomputer. We attribute the relatively low impact of omitting accelerations to the strictly cyclic nature and low variability of walking at a constant speed on a treadmill. Accelerations remain of critical importance for data-driven control when extending to diverse tasks, which is supported by their theoretical relevance and incorporation by related work [13, 12]. Nevertheless, research has explored gait event detection from solely gyroscope data [24, 25], and our results utilizing angular information may warrant further exploration for sensor minimization in strictly cyclic tasks.
@@ III-B Optimal Model
The results indicate that TCN-based ankle moment estimation is promising in the context of existing literature. [26] estimated post-stroke ankle dorsi-plantarflexion torque using an EMG-driven musculoskeletal model combined with data from five IMUs. They reported leave-one-out accuracies of $\mathrm{R}^{2}$ 0.67±0.220.67$\pm$ 0.220.67 ± 0.22 and MAE 0.32±0.170.32$\pm$ 0.170.32 ± 0.17 Nm/kg at comfortable walking speeds (CWS) and $\mathrm{R}^{2}$ 0.63±0.200.63$\pm$ 0.200.63 ± 0.20 and MAE 0.39±0.150.39$\pm$ 0.150.39 ± 0.15 Nm/kg at fast walking speeds (FWS). Our TCN models achieved comparable or improved performance while relying on only three IMUs placed on the paretic limb. However, this comparison serves only as an indication, as we validated our model on a different dataset. Similarly, [20] estimated ankle torques in healthy individuals walking overground using EMGs and seven IMUs. They reported leave-one-out errors of MAE 0.158±0.0870.158$\pm$ 0.0870.158 ± 0.087 Nm/kg ($\mathrm{R}^{2}$=74) when using a CNN and MAE 0.117±0.0830.117$\pm$ 0.0830.117 ± 0.083 Nm/kg ($\mathrm{R}^{2}$=84) with a Random Forest model. Although their results were obtained on a different population and sensor setup, our work suggests that data-driven methods using fewer sensors can still provide accurate torque estimates for post-stroke populations.
Fine-tuning a model pretrained on the Unimpaired Dataset had an observable positive effect on leave-one-out model performance for TCN2OUT. Without pretraining, TCN2OUT achieved $\mathrm{R}^{2}$ 0.65±0.160.65$\pm$ 0.160.65 ± 0.16, whereas pretraining increased this to $\mathrm{R}^{2}$ 0.74±0.130.74$\pm$ 0.130.74 ± 0.13, which is the highest among all compared models. This suggests that pretraining helped TCN2OUT optimize both ankle moment estimation and stance phase classification, supporting our hypothesis that learning both tasks may have encouraged the model to develop a more structured, physiologically meaningful representation of gait dynamics.
@@ III-C Exoskeleton Application
Testing our prototype, we collected 159 strides at comfortable speed from two walking bouts. We utilized the TCN2OUT for ankle torque estimates, which achieved an overall RMSE of 0.14 Nm/kg, MAE 0.1 Nm/kg, and $\mathrm{R}^{2}$ of 0.63. These results are comparable to results on our Post-stroke Dataset (Table II). We evaluated observed dorsiflexion (DF) and plantarflexion (PF) peaks timing in % gait cycle and magnitude. These measurements are of high importance for exoskeleton controllers as they decide when and how much assistance is applied. As depicted in Figure 5, these peaks occur around initial contact and push-off in the gait cycle. The results are summarized in Figure 6. The PF peaks exhibited a notable error of 0.24±0.090.24$\pm$ 0.090.24 ± 0.09 Nm/kg. The timing error of both peaks however was promisingly low with 1.94±0.971.94$\pm$ 0.971.94 ± 0.97 % gait cycle deviation for DF peaks and 1.56±0.951.56$\pm$ 0.951.56 ± 0.95 % gait cycle for PF peaks. Given the low error in timing, the higher magnitude error was not a major concern for our exosuit trial, as it can be adjusted through proportional gain by the controller. Considering the relatively small population of four post-stroke participants in our training data, we expect that training future estimation models on larger populations could improve these results further.
@@
Our evaluation involved ankle torque estimates for the entire stride, including swing-phase. Although swing phase predictions carry limited significance due to the relatively low ankle torque the paretic limb produces during swing, they remain critical for ankle dorsiflexion and toe clearance. As observed in Table II, TCN2OUT achieved lower RMSEs for swing phase predictions (0.05±0.0340.05$\pm$ 0.0340.05 ± 0.034 Nm/kg for fine-tuned models, 0.08±0.0570.08$\pm$ 0.0570.08 ± 0.057 Nm/kg for post-stroke-only models). However, the ground-truth data used for training was not fully representative of true swing-phase ankle torque due to limitations in modeling the inertial properties of the body segments used to compute inverse dynamics. Future work could address this by incorporating the stance phase predictions of a dual-output model like TCN2OUT to affect how the controller utilizes the ankle torque estimate.

### v2

@@ II-C1 Unimpaired Dataset
The Unimpaired Dataset was collected from individuals without neuromotor gait impairment (6 participants, 4 M / 2 F, age $29\pm 3.9$ yrs). We sensorized their left leg, as all participants post-stroke were left-paretic. As shown in Figure 3 on the left, all the IMUs were laterally positioned on the participants’ left leg, aligned with the sagittal plane. Participants walked on the instrumented treadmill for 6 trials of 4 minutes each. Each trial featured randomized speed conditions ranging from 0.6 m/s to 1.4 m/s. For the first 2 minutes, the treadmill maintained a constant speed of 1 m/s before ramping up to either 1.2 m/s or 1.4 m/s, or down to 0.8 m/s or 0.6 m/s for the final 2 minutes. Half of the participants started with acceleration, while the others began with deceleration. The speed variations were implemented to capture a range of walking patterns and enhance dataset diversity. On average, $2118\pm 442$ strides were collected per participant, resulting in a total of $12710$ strides.
@@ II-C2 Post-stroke Dataset
The Post-stroke Dataset was collected from 4 left-paretic stroke survivors (4 M, age $55\pm 5.5$ yrs). Participants wore the ankle robot on their paretic limb. Each participant completed 6 trials on the instrumented treadmill, comprising of two 2 min trials for each a self-selected comfortable walking speed (CWS: $0.65\pm 0.13$ m/s), a slow walking speed (SWS: $0.5\pm 0.08$ m/s), and a fast walking speed (FWS: $0.8\pm 0.18$ m/s). Participants post-stroke demonstrated high variability in their self-selected CWS, which ranged from 0.5 to 0.8 m/s. On average, each participant performed $334\pm 79$ strides, which resulted in a total of $1334$ strides.
@@ III-A Feature Importance
We conducted a feature importance analysis to reduce the number of streamed channels. The results are summarized in Table I. Randomizing the acceleration measurements increased RMSE by $0.04\pm 0.03$ Nm/kg, the MAE by $0.03\pm 0.02$ Nm/kg, and decreased the $\mathrm{R}^{2}$ score by $0.09\pm 0.05$ on average. Despite accelerations being integral to inverse dynamics, their omission affected prediction accuracy the least, compared to orientation and gyroscope data. We excluded the accelerations from our subsequent model training to ensure the targeted 100 Hz throughput on the microcomputer. We attribute the relatively low impact of omitting accelerations to the strictly cyclic nature and low variability of walking at a constant speed on a treadmill. Accelerations remain of critical importance for data-driven control when extending to diverse tasks, which is supported by their theoretical relevance and incorporation by related work [13, 12]. Nevertheless, research has explored gait event detection from solely gyroscope data [24, 25], and our results utilizing angular information may warrant further exploration for sensor minimization in strictly cyclic tasks.
@@ III-B Optimal Model
The results indicate that TCN-based ankle moment estimation is promising in the context of existing literature. [26] estimated post-stroke ankle dorsi-plantarflexion torque using an EMG-driven musculoskeletal model combined with data from five IMUs. They reported leave-one-out accuracies of $\mathrm{R}^{2}$ $0.67\pm 0.22$ and MAE $0.32\pm 0.17$ Nm/kg at comfortable walking speeds (CWS) and $\mathrm{R}^{2}$ $0.63\pm 0.20$ and MAE $0.39\pm 0.15$ Nm/kg at fast walking speeds (FWS). Our TCN models achieved comparable or improved performance while relying on only three IMUs placed on the paretic limb. However, this comparison serves only as an indication, as we validated our model on a different dataset. Similarly, [20] estimated ankle torques in healthy individuals walking overground using EMGs and seven IMUs. They reported leave-one-out errors of MAE $0.158\pm 0.087$ Nm/kg ($\mathrm{R}^{2}$=74) when using a CNN and MAE $0.117\pm 0.083$ Nm/kg ($\mathrm{R}^{2}$=84) with a Random Forest model. Although their results were obtained on a different population and sensor setup, our work suggests that data-driven methods using fewer sensors can still provide accurate torque estimates for post-stroke populations.
Fine-tuning a model pretrained on the Unimpaired Dataset had an observable positive effect on leave-one-out model performance for TCN2OUT. Without pretraining, TCN2OUT achieved $\mathrm{R}^{2}$ $0.65\pm 0.16$, whereas pretraining increased this to $\mathrm{R}^{2}$ $0.74\pm 0.13$, which is the highest among all compared models. This suggests that pretraining helped TCN2OUT optimize both ankle moment estimation and stance phase classification, supporting our hypothesis that learning both tasks may have encouraged the model to develop a more structured, physiologically meaningful representation of gait dynamics.
@@ III-C Exoskeleton Application
Testing our prototype, we collected 159 strides at comfortable speed from two walking bouts. We utilized the TCN2OUT for ankle torque estimates, which achieved an overall RMSE of 0.14 Nm/kg, MAE 0.1 Nm/kg, and $\mathrm{R}^{2}$ of 0.63. These results are comparable to results on our Post-stroke Dataset (Table II). We evaluated observed dorsiflexion (DF) and plantarflexion (PF) peaks timing in % gait cycle and magnitude. These measurements are of high importance for exoskeleton controllers as they decide when and how much assistance is applied. As depicted in Figure 5, these peaks occur around initial contact and push-off in the gait cycle. The results are summarized in Figure 6. The PF peaks exhibited a notable error of $0.24\pm 0.09$ Nm/kg. The timing error of both peaks however was promisingly low with $1.94\pm 0.97$ % gait cycle deviation for DF peaks and $1.56\pm 0.95$ % gait cycle for PF peaks. Given the low error in timing, the higher magnitude error was not a major concern for our exosuit trial, as it can be adjusted through proportional gain by the controller. Considering the relatively small population of four post-stroke participants in our training data, we expect that training future estimation models on larger populations could improve these results further.
@@
Our evaluation involved ankle torque estimates for the entire stride, including swing-phase. Although swing phase predictions carry limited significance due to the relatively low ankle torque the paretic limb produces during swing, they remain critical for ankle dorsiflexion and toe clearance. As observed in Table II, TCN2OUT achieved lower RMSEs for swing phase predictions ($0.05\pm 0.034$ Nm/kg for fine-tuned models, $0.08\pm 0.057$ Nm/kg for post-stroke-only models). However, the ground-truth data used for training was not fully representative of true swing-phase ankle torque due to limitations in modeling the inertial properties of the body segments used to compute inverse dynamics. Future work could address this by incorporating the stance phase predictions of a dual-output model like TCN2OUT to affect how the controller utilizes the ankle torque estimate.